# Revision of conservative lower bound on intergalactic magnetic field from Fermi and Cherenkov telescope observations of extreme blazars

### v1

@@ 3.2 Sources that yield constraints on IGMF
Following Aharonian et al. (2023), we also consider the combination of IGMF lower bounds by adding the $\Delta\chi^{2}$ for all the seven sources. Similar to the single source analysis, for each value of $B$, we find the best fit intrinsic spectral parameters $E_{cut,i},\Gamma_{i},N_{i}$ for each $i$th source separately, but the value of $B$ is common for all sources. Adopting the approach of Avni (1976) also for this case, we assume that the cumulative $\chi^{2}(B|E_{cut,1},\Gamma_{cut,1},N_{1},...E_{cut,7},\Gamma_{7},N_{7})$ also follows the $\chi^{2}$ statistics with one degree of freedom. Combination of data on all sources slightly relaxes the lower bound on IGMF, down to
$B_{min,all}>1.3\times 10^{-17}\mbox{ G.}$. (3)
@@ 4 Discussion and conclusions
We have used the set of extreme blazars detected in the VHE band by both Fermi/LAT and IACTs (Neronov & Semikoz, 2025) to revise the lower bound on magnetic fields in the voids of the Large Scale Structure. This bound stems from the observation of the effect of suppression of the secondary $\gamma$-ray flux from electron-positron pairs deposited in the voids in result of interactions of VHE $\gamma$-rays with the optical and infrared photons of the EBL. Similar to Acciari et al. (2023a); Aharonian et al. (2023), we consider ”conservative” (perhaps ”over-conservative”) assumption that sources have only ”switched on” at the start of observations in the VHE band and the secondary emission is only suppressed because of the time delay of secondary flux caused by deflections of electrons and positrons in the IGMF.
@@
In the specific case of the blazar 1ES 0229+200, considered by Acciari et al. (2023a), our bound is somewhat weaker than that derived by Acciari et al. (2023a), because we do not follow the details of source variability over the 20 years of source observations in the VHE band and assume that the source luminosity is represented by a $\theta$-function like switch-on followed by the constant luminosity on the 16-year time scale (representative of the time-averaged source luminosity). Comparing our limit on IGMF with that derived by Aharonian et al. (2023), we find that we get a weaker lower bound, even though our modeling of source ”switch-on” is the same as that of Aharonian et al. (2023). We argue that the origin of the mismatch of our and Acciari et al. (2023a) results with the results of Aharonian et al. (2023) stems from a discrepancy in the modelling of the secondary flux (see Appendix A for details).
@@
Even though our limit on IGMF from the data on 1ES 0229+200 is weaker than that of Acciari et al. (2023a), we have identified another source, 1ES 0502+675, that provides a tighter lower bound on IGMF, at the level of $B>2.1\times 10^{-17}$ G. We have checked that combination of the data on all sources that yield constraints on IGMF (seven in total) slightly worsens the lower bound on IGMF down to $B>1.3\times 10^{-17}$ G. It is surprising that 1ES 0502+675 has not been considered before in the IGMF search and even no refereed journal publication [^1] A powerlaw spectral fit for this source is given in Abeysekara et al. (2019). We show this fit in Fig. 1. about the source spectrum has appeared since the detection of the source back in 2009 by VERITAS (Ong, 2009). We were only able to find a conference proceedings report on the source spectrum. Given that the source provides the tightest constraints on IGMF in our analysis, we believe that a better characterization of its TeV band spectrum with longer IACT exposures would be useful. The source is very bright (the second-brightest VHE extreme blazar after Mrk 501, as noticed in Neronov & Semikoz (2025)), so that detailed characterization of its spectrum in the TeV band should be possible already with the current generation IACTs. There is clearly a potential for further tightening of the conservative lower bound on IGMF via extension of the IACT measurements in the energy range above 1 TeV.
@@
In our analysis we have focused on IGMF with large correlation length $\lambda_{B}>l_{e}$. Nevertheless, we can extrapolate the results to the case of shows correlation length fields using prescriptions of Neronov & Semikoz (2009). The lower bound is independent of $\lambda_{B}$ as long as $\lambda_{B}>l_{e}$ and it scales as $\lambda_{B}^{-1/2}$ in the $\lambda_{B}<l_{e}$ range. The estimate of $l_{e}$ in our analysis is identical to that of Acciari et al. (2023a), $l_{e}\sim 0.1$ Mpc, so that our bound for the short correlation length field is
@@
The short correlation length field bound is particularly interesting for the cosmological IGMF that may originate from cosmological phase transitions (Electroweak, quark confinement) (Durrer & Neronov, 2013). For such fields the correlation length and strength are related by the ”largest processed eddy” condition. There exist alternative estimates of the largest processed eddy scale, with the estimate of Banerjee & Jedamzik (2004) is $\lambda\sim 1[B/10^{-8.5}\mbox{ G}]$ Mpc. Hosking & Schekochihin (2023); Brandenburg et al. (2024) estimate it to be two-to-four orders of magnitude shorter. This gives a lower bound of $\sim 10^{-14}$ G on the cosmological magnetic field under the assumption of Banerjee & Jedamzik (2004) largest processed eddy scaling and up to two orders of magnitude stronger assuming Hosking & Schekochihin (2023) largest processed eddy scaling. The cosmological relic magnetic field with the strength close to this lower bound can, for example, be helical magnetic field that was responsible for the generation of baryon asymmetry of the Universe during the Electroweak cross-over (Kamada & Long, 2016; Boyer & Neronov, 2025), or it can be a non-helical field generated at quark confinement first order phase transition (Sigl et al., 1997).
The field at the level close to the lower bound derived above should be detectable using multiple techniques: through detection of delayed ”pair echo” emission in the signals of AGN (Acciari et al., 2023a) and Gamma-ray Bursts (Miceli et al., 2024; Takahashi et al., 2011), or through the detection of extended emission around AGN (Neronov & Semikoz, 2009; Ackermann et al., 2018; Aharonian et al., 2023). To identify best targets for such search, it is necessary to improve the knowledge of the intrinsic spectra of sources from Table 1: for most of them, no certain predictions for the secondary signal flux can be made because of uncetainty of the high-energy cut-offs in the spectra. This can be improved using the new generation of IACTs: CTAO (Cherenkov Telescope Array Consortium et al., 2019) and LACT (Zhang et al., 2025) that will have better sensitivity in the multi-TeV energy range. The information on the time-average intrinsic fluxes of the sources and on the presence of high-energy cut-offs in their spectra can also be provided by air shower arrays, like already operating LHAASO (Cao et al., 2024) and planned SWGO (Huentemeyer et al., 2019).
@@ Acknowledgements.
We would like to thank J.Biteau for clarifying discussions. This work has been supported in part by the French National Research Agency (ANR) grant ANR-24-CE31-4686.
@@
Appendix A Modeling of the secondary flux and comparison with Aharonian et al. (2023)
@@
Figure 5: Comparison of the secondary flux from Aharonian et al. (2023) with our secondary flux model for the same parameters (magnetic field $B=3.2\times 10^{-16}$ G, activity time $T=10$ yr, source 1ES 0229+200). Orange data points show the result of Fermi/LAT likelihood analysis by Aharonian et al. (2023).
@@
We compute the secondary flux using the CRbeam code (Berezinsky & Kalashev 2016; Kalashev et al. 2023). We simulate a beam of photons in the energy range between 10 GeV and 30 TeV, with $dN/dE\propto E^{-1}$ spectrum. For each secondary photon produced following the $e^{+}e^{-}$ pair production followed by the inverse Compton scattering, we keep a memory of the primary photon. This allows us to calculate the secondary flux for any primary photon spectrum using single simulation of the $E^{-1}$ spectrum, by simply weighting the secondary photons according to the energy of the primary photon. We propagate primary and secondary particles until they cross a sphere of the radius $D$ equal to the distance to the source of interest. For each photon crossing the sphere, we calculate the angle between the direction of the photon and normal to the sphere, which corresponds to the angle of misalignment of the arrival direction of the secondary photons. We also notice the total path from the source for each photon crossing the sphere. This path, divided by the speed of light, gives the arrival time of all photons.
@@
Fig. 5 shows a comparison of our calculation of the secondary flux for $T=10$ yr activity time of 1ES 0229+200, for the magnetic field $B=3.2\times 10^{-16}$ G with that of Aharonian et al. (2023). Our estimate of the secondary flux is by up to a factor-of-two lower than that of Aharonian et al. (2023). This difference does not stem from the discrepancies between CRbeam code that we use and CRpropa code (Alves Batista et al. 2022) used by Aharonian et al. (2023), because the two codes have been found to agree within the redshift range of interest (Kalashev et al. 2023). CRpropa code was also used in the modeling of the secondary flux by Acciari et al. (2023b), which which our calculations agree.
@@
Our analysis is based on Fermi/LAT spectra extracted using the aperture photometry technique, as compared to the likelihood analysis used by Aharonian et al. (2023). To estimate the source flux in each energy bin, we use the knowledge of the 68% and 95% containment radii of the Fermi/LAT point spread function. If the total source signal in the energy bin is $S$ and the background per square degree is $b$, the numbers of photon counts in the 68% and 95% containment circles are
@@
We use this estimate in the energy range below 50 GeV, where the background level is high. In the energy range above 50 GeV, the sources are mostly detected in the background-free regime and we assume that $S=S_{95}/0.95$. The aperture photometry approach has an advantage compared to the likelihood analysis approach in the ”uncrowded” sky regions where there are no sources with overlapping signals. Contrary to the likelihood estimate, it does not depend on the sky model that has many parameters (of the Galactic diffuse emission, isotropic diffuse emission, distribution of point sources in the region of interest, choice of the spectral models and spectral parameters of those sources). We have checked that our aperture photometry spectra are consistent with those of the likelihood analysis by Aharonian et al. (2023). Fig. 5 shows a comparison of the aperture photometry and likelihood analysis for 1ES 0229+200.

### v2

@@ 3.2 Sources that yield constraints on IGMF
Following Aharonian and others (2023), we also consider the combination of IGMF lower bounds by adding the $\Delta\chi^{2}$ for all the seven sources. Similar to the single source analysis, for each value of $B$, we find the best fit intrinsic spectral parameters $E_{cut,i},\Gamma_{i},N_{i}$ for each $i$th source separately, but the value of $B$ is common for all sources. Adopting the approach of Avni (1976) also for this case, we assume that the cumulative $\chi^{2}(B|E_{cut,1},\Gamma_{cut,1},N_{1},...E_{cut,7},\Gamma_{7},N_{7})$ also follows the $\chi^{2}$ statistics with one degree of freedom. Combination of data on all sources slightly relaxes the lower bound on IGMF, down to
$B_{min,all}>1.3\times 10^{-17}\mbox{ G.}$ (3)
@@ 4 Discussion and conclusions
We have used the set of extreme blazars detected in the VHE band by both Fermi/LAT and IACTs (Neronov and Semikoz, 2025) to revise the lower bound on magnetic fields in the voids of the Large Scale Structure. This bound stems from the observation of the effect of suppression of the secondary $\gamma$-ray flux from electron-positron pairs deposited in the voids in result of interactions of VHE $\gamma$-rays with the optical and infrared photons of the EBL. Similar to Acciari et al. (2023a); Aharonian and others (2023), we consider ”conservative” (perhaps ”over-conservative”) assumption that sources have only ”switched on” at the start of observations in the VHE band and the secondary emission is only suppressed because of the time delay of secondary flux caused by deflections of electrons and positrons in the IGMF.
@@
In the specific case of the blazar 1ES 0229+200, considered by Acciari et al. (2023a), our bound is somewhat weaker than that derived by Acciari et al. (2023a), because we do not follow the details of source variability over the 20 years of source observations in the VHE band and assume that the source luminosity is represented by a $\theta$-function like switch-on followed by the constant luminosity on the 16-year time scale (representative of the time-averaged source luminosity). Comparing our limit on IGMF with that derived by Aharonian and others (2023), we find that we get a weaker lower bound, even though our modeling of source ”switch-on” is the same as that of Aharonian and others (2023). We argue that the origin of the mismatch of our and Acciari et al. (2023a) results with the results of Aharonian and others (2023) stems from a discrepancy in the modelling of the secondary flux (see Appendix A for details).
@@
Even though our limit on IGMF from the data on 1ES 0229+200 is weaker than that of Acciari et al. (2023a), we have identified another source, 1ES 0502+675, that provides a tighter lower bound on IGMF, at the level of $B>2.1\times 10^{-17}$ G. We have checked that combination of the data on all sources that yield constraints on IGMF (seven in total) slightly worsens the lower bound on IGMF down to $B>1.3\times 10^{-17}$ G. It is surprising that 1ES 0502+675 has not been considered before in the IGMF search and even no refereed journal publication [^1] A powerlaw spectral fit for this source is given in Abeysekara et al. (2019). We show this fit in Fig. 1. about the source spectrum has appeared since the detection of the source back in 2009 by VERITAS (Ong, 2009). We were only able to find a conference proceedings report on the source spectrum. Given that the source provides the tightest constraints on IGMF in our analysis, we believe that a better characterization of its TeV band spectrum with longer IACT exposures would be useful. The source is very bright (the second-brightest VHE extreme blazar after Mrk 501, as noticed in Neronov and Semikoz (2025)), so that detailed characterization of its spectrum in the TeV band should be possible already with the current generation IACTs. Such measurements could better constrain the energy cut-off and possibly extend the spectrum in the energy range above 1 TeV, which would allow for a further tightening of the conservative lower bound on IGMF.
A potential caveat to these constraints is the relaxation of the electron-positron pair beam via collective plasma instabilities, which could dissipate the beam energy into the intergalactic medium rather than producing secondary $\gamma$-rays. Several studies have argued that the cooling time due to interactions between the relativistic pair beams and the intergalactic medium can be shorter than the inverse-Compton cooling time translating to plasma instabilities that can reduce the secondary cascade flux (Broderick et al., 2012, 2018; Vafin et al., 2019) and lead to smaller inferred lower bound on the IGMF of blazars (e.g. see Dey and Sigl (2025) for the case of 1ES 0229+200). Some models suggest that large-scale magnetic fields might suppress the growth of these instabilities (Alawashra and Pohl, 2022; Alawashra et al., 2025). The results of the laboratory experiment in Arrowsmith and others (2025) demonstrated that the instability growth rates are heavily suppressed unless the particle beam is perfectly collimated or monochromatic. These ideal conditions are usually not met by blazar-induced pairs. Hence, inferring lower limits on the IGMF strength from non-observation of GeV cascade of blazars remains a robust method.
@@
In our analysis we have focused on IGMF with large correlation length $\lambda_{B}>l_{e}$. Nevertheless, we can extrapolate the results to the case of shows correlation length fields using prescriptions of Neronov and Semikoz (2009). The lower bound is independent of $\lambda_{B}$ as long as $\lambda_{B}>l_{e}$ and it scales as $\lambda_{B}^{-1/2}$ in the $\lambda_{B}<l_{e}$ range. The estimate of $l_{e}$ in our analysis is identical to that of Acciari et al. (2023a), $l_{e}\sim 0.1$ Mpc, so that our bound for the short correlation length field is
@@
The short correlation length field bound is particularly interesting for the cosmological IGMF that may originate from cosmological phase transitions (Electroweak, quark confinement) (Durrer and Neronov, 2013). For such fields the correlation length and strength are related by the ”largest processed eddy” condition. There exist alternative estimates of the largest processed eddy scale, with the estimate of Banerjee and Jedamzik (2004) is $\lambda\sim 1[B/10^{-8.5}\mbox{ G}]$ Mpc. Hosking and Schekochihin (2023); Brandenburg et al. (2024) estimate it to be two-to-four orders of magnitude shorter. This gives a lower bound of $\sim 10^{-14}$ G on the cosmological magnetic field under the assumption of Banerjee and Jedamzik (2004) largest processed eddy scaling and up to two orders of magnitude stronger assuming Hosking and Schekochihin (2023) largest processed eddy scaling. The cosmological relic magnetic field with the strength close to this lower bound can, for example, be helical magnetic field that was responsible for the generation of baryon asymmetry of the Universe during the Electroweak cross-over (Kamada and Long, 2016; Boyer and Neronov, 2025), or it can be a non-helical field generated at quark confinement first order phase transition (Sigl et al., 1997).
The field at the level close to the lower bound derived above should be detectable using multiple techniques: through detection of delayed ”pair echo” emission in the signals of AGN (Acciari et al., 2023a) and Gamma-ray Bursts (Miceli et al., 2024; Takahashi et al., 2011), or through the detection of extended emission around AGN (Neronov and Semikoz, 2009; Ackermann et al., 2018; Aharonian and others, 2023). To identify best targets for such search, it is necessary to improve the knowledge of the intrinsic spectra of sources from Table 1: for most of them, no certain predictions for the secondary signal flux can be made because of uncetainty of the high-energy cut-offs in the spectra. This can be improved using the new generation of IACTs: CTAO (Cherenkov Telescope Array Consortium et al., 2019) and LACT (Zhang et al., 2025) that will have better sensitivity in the multi-TeV energy range. The information on the time-average intrinsic fluxes of the sources and on the presence of high-energy cut-offs in their spectra can also be provided by air shower arrays, like already operating LHAASO (Cao et al., 2024) and planned SWGO (Huentemeyer et al., 2019).
@@ Acknowledgements.
We would like to thank J. Biteau for clarifying discussions as well as the anonymous referee for their useful suggestions to improve the manuscript. This work has been supported in part by the French National Research Agency (ANR) grant ANR-24-CE31-4686.
@@
Appendix A Modeling of the secondary flux and comparison with Aharonian and others (2023)
@@
Figure 5: Comparison of the secondary flux from Aharonian and others (2023) with our secondary flux model for the same parameters (magnetic field $B=3.2\times 10^{-16}$ G, activity time $T=10$ yr, source 1ES 0229+200). Results from CRbeam (black dotted line) and CRpropa (red dotted line) are both present. Orange data points show the result of Fermi/LAT likelihood analysis by Aharonian and others (2023).
@@
We compute the secondary flux using the CRbeam code (Berezinsky and Kalashev 2016; Kalashev et al. 2023). We simulate a beam of photons in the energy range between 10 GeV and 30 TeV, with $dN/dE\propto E^{-1}$ spectrum. For each secondary photon produced following the $e^{+}e^{-}$ pair production followed by the inverse Compton scattering, we keep a memory of the primary photon. This allows us to calculate the secondary flux for any primary photon spectrum using single simulation of the $E^{-1}$ spectrum, by simply weighting the secondary photons according to the energy of the primary photon. We propagate primary and secondary particles until they cross a sphere of the radius $D$ equal to the distance to the source of interest. For each photon crossing the sphere, we calculate the angle between the direction of the photon and normal to the sphere, which corresponds to the angle of misalignment of the arrival direction of the secondary photons. We also notice the total path from the source for each photon crossing the sphere. This path, divided by the speed of light, gives the arrival time of all photons. For completeness, we run the same configuration considered in CRbeam with CRpropa (Alves Batista et al. 2022) and calculate the corresponding secondary flux following the procedure described above.
@@
Fig. 5 shows a comparison of our calculation of the secondary flux for $T=10$ yr activity time of 1ES 0229+200, for the magnetic field $B=3.2\times 10^{-16}$ G with that of Aharonian and others (2023) [^2] The original secondary flux curve has been extracted from slide 28 of the following presentation: https://indico.ict.inaf.it/event/3156/contributions/21372/attachments/9760/20166/MMeyer_IGMF_Trieste_2025.pdf. Our estimate of the secondary flux is by up to a factor-of-two lower than that of Aharonian and others (2023). This difference does not stem from the discrepancies between CRbeam code and CRpropa code used by Aharonian and others (2023) as shown by Fig. A.1, which is consistent with Kalashev et al. (2023) where the two codes have been found to agree within the redshift range of interest. CRpropa code was also used in the modeling of the secondary flux by Acciari et al. (2023b), with which our calculations agree.
@@
Our analysis is based on Fermi/LAT spectra extracted using the aperture photometry technique, as compared to the likelihood analysis used by Aharonian and others (2023). To estimate the source flux in each energy bin, we use the knowledge of the 68% and 95% containment radii of the Fermi/LAT point spread function. If the total source signal in the energy bin is $S$ and the background per square degree is $b$, the numbers of photon counts in the 68% and 95% containment circles are
@@
We use this estimate in the energy range below 50 GeV, where the background level is high. In the energy range above 50 GeV, the sources are mostly detected in the background-free regime and we assume that $S=S_{95}/0.95$. The aperture photometry approach has an advantage compared to the likelihood analysis approach in the ”uncrowded” sky regions where there are no sources with overlapping signals. Contrary to the likelihood estimate, it does not depend on the sky model that has many parameters (of the Galactic diffuse emission, isotropic diffuse emission, distribution of point sources in the region of interest, choice of the spectral models and spectral parameters of those sources). We have checked that our aperture photometry spectra are consistent with those of the likelihood analysis by Aharonian and others (2023). Fig. 5 shows a comparison of the aperture photometry and likelihood analysis for 1ES 0229+200.
Table 3: Best-fit spectral parameters of the sources listed in Table 2 for different IGMF strengths.
$B$ [G] $N_{0}$ $10^{-12}$ TeV-1cm-2s-1] $\Gamma$ $E_{cut}$ [TeV]
1ES 0229+200
$6.00\times 10^{-19}$ $3.63^{+0.74}_{-0.61}$ $1.17^{+0.12}_{-0.16}$ $1.78^{+0.41}_{-0.30}$
$1.20\times 10^{-18}$ $3.55^{+0.72}_{-0.53}$ $1.17^{+0.12}_{-0.17}$ $1.86^{+0.43}_{-0.35}$
$2.39\times 10^{-18}$ $3.47^{+0.60}_{-0.52}$ $1.16^{+0.13}_{-0.16}$ $2.09^{+0.48}_{-0.39}$
$4.76\times 10^{-18}$ $3.24^{+0.56}_{-0.49}$ $1.18^{+0.13}_{-0.17}$ $2.45^{+0.64}_{-0.45}$
$9.50\times 10^{-18}$ $2.95^{+0.52}_{-0.44}$ $1.25^{+0.12}_{-0.15}$ $3.16^{+0.91}_{-0.65}$
$1.89\times 10^{-17}$ $2.69^{+0.47}_{-0.40}$ $1.35^{+0.10}_{-0.13}$ $4.27^{+1.62}_{-1.03}$
$3.78\times 10^{-17}$ $2.51^{+0.44}_{-0.42}$ $1.44^{+0.08}_{-0.11}$ $5.62^{+3.29}_{-1.73}$
$7.54\times 10^{-17}$ $2.34^{+0.41}_{-0.43}$ $1.52^{+0.06}_{-0.08}$ $7.41^{+6.08}_{-2.73}$
$1.50\times 10^{-16}$ $2.45^{+0.50}_{-0.54}$ $1.56^{+0.06}_{-0.06}$ $7.24^{+8.98}_{-2.77}$
$3.00\times 10^{-16}$ $2.19^{+0.50}_{-0.64}$ $1.60^{+0.07}_{-0.06}$ $10.72^{+20.90}_{-5.35}$
1ES 0347-121
$6.00\times 10^{-19}$ $7.08^{+1.63}_{-1.33}$ $1.36^{+0.07}_{-0.08}$ $0.83^{+0.22}_{-0.17}$
$1.20\times 10^{-18}$ $6.46^{+1.67}_{-1.21}$ $1.39^{+0.07}_{-0.08}$ $0.91^{+0.26}_{-0.20}$
$2.39\times 10^{-18}$ $5.75^{+1.33}_{-1.08}$ $40^{+0.07}_{-0.08} 1.1$$12^{+0.36}_{-0.27}$
$4.76\times 10^{-18}$ $5.13^{+1.33}_{-0.96}$ $1.45^{+0.06}_{-0.07}$ $1.29^{+0.53}_{-0.38}$
$9.50\times 10^{-18}$ $4.68^{+1.21}_{-0.88}$ $1.50^{+0.06}_{-0.06}$ $1.48^{+0.76}_{-0.46}$
$1.89\times 10^{-17}$ $4.47^{+1.29}_{-0.92}$ $1.54^{+0.05}_{-0.06}$ $1.58^{+0.99}_{-0.54}$
$3.78\times 10^{-17}$ $4.37^{+1.26}_{-0.98}$ $1.57^{+0.05}_{-0.06}$ $1.55^{+1.21}_{-0.55}$
$7.54\times 10^{-17}$ $3.98^{+1.15}_{-0.89}$ $1.59^{+0.06}_{-0.06}$ $1.95^{+2.12}_{-0.77}$
$1.50\times 10^{-16}$ $3.89^{+1.24}_{-0.94}$ $1.60^{+0.06}_{-0.07}$ $2.04^{+2.13}_{-0.92}$
$3.00\times 10^{-16}$ $3.80^{+1.21}_{-0.92}$ $1.61^{+0.07}_{-0.07}$ $2.24^{+2.55}_{-1.01}$
1ES 0502+675
$6.00\times 10^{-19}$ $93.33^{+16.32}_{-13.90}$ $1.23^{+0.04}_{-0.04}$ $0.35^{+0.03}_{-0.03}$
$1.20\times 10^{-18}$ $91.20^{+15.95}_{-13.58}$ $1.23^{+0.04}_{-0.04}$ $0.38^{+0.04}_{-0.03}$
$2.39\times 10^{-18}$ $89.13^{+15.58}_{-13.27}$ $1.23^{+0.04}_{-0.04}$ $0.43^{+0.04}_{-0.04}$
$4.76\times 10^{-18}$ $83.18^{+14.54}_{-12.39}$ $1.23^{+0.04}_{-0.05}$ $0.52^{+0.06}_{-0.06}$
$9.50\times 10^{-18}$ $75.86^{+13.27}_{-11.30}$ $1.22^{+0.05}_{-0.05}$ $0.68^{+0.08}_{-0.07}$
$1.89\times 10^{-17}$ $63.10^{+9.34}_{-9.40}$ $1.23^{+0.06}_{-0.06}$ $0.98^{+0.14}_{-0.15}$
$3.78\times 10^{-17}$ $56.23^{+9.83}_{-8.37}$ $1.19^{+0.09}_{-0.11}$ $1.55^{+0.23}_{-0.23}$
$7.54\times 10^{-17}$ $44.67^{+6.62}_{-6.65}$ $1.32^{+0.05}_{-0.09}$ $2.09^{+0.54}_{-0.50}$
$1.50\times 10^{-16}$ $39.81^{+6.96}_{-5.93}$ $1.40^{+0.04}_{-0.04}$ $2.29^{+0.80}_{-0.67}$
$3.00\times 10^{-16}$ $37.15^{+6.50}_{-6.25}$ $1.44^{+0.03}_{-0.03}$ $2.63^{+1.35}_{-0.93}$
1ES 1101-232
$6.00\times 10^{-19}$ $10.00^{+2.02}_{-1.49}$ $1.20^{+0.09}_{-0.11}$ $0.87^{+0.13}_{-0.11}$
$1.20\times 10^{-18}$ $9.12^{+1.84}_{-1.36}$ $1.22^{+0.08}_{-0.10}$ $0.98^{+0.17}_{-0.15}$
$2.39\times 10^{-18}$ $8.71^{+1.52}_{-1.30}$ $22^{+0.09}_{-0.10} 1.1$$12^{+0.20}_{-0.17}$
$4.76\times 10^{-18}$ $7.94^{+1.39}_{-1.18}$ $1.23^{+0.09}_{-0.11}$ $1.35^{+0.24}_{-0.20}$
$9.50\times 10^{-18}$ $6.92^{+1.02}_{-0.89}$ $1.28^{+0.09}_{-0.11}$ $1.74^{+0.35}_{-0.29}$
$1.89\times 10^{-17}$ $6.03^{+0.89}_{-0.78}$ $1.35^{+0.08}_{-0.10}$ $2.29^{+0.59}_{-0.47}$
$3.78\times 10^{-17}$ $5.37^{+0.80}_{-0.80}$ $1.45^{+0.06}_{-0.07}$ $2.82^{+1.26}_{-0.78}$
$7.54\times 10^{-17}$ $4.79^{+0.84}_{-0.81}$ $1.50^{+0.05}_{-0.06}$ $3.72^{+2.59}_{-1.26}$
$1.50\times 10^{-16}$ $4.47^{+0.78}_{-1.00}$ $1.54^{+0.05}_{-0.05}$ $4.47^{+5.08}_{-1.78}$
$3.00\times 10^{-16}$ $4.27^{+0.86}_{-1.51}$ $1.57^{+0.08}_{-0.05}$ $5.37^{+16.01}_{-2.42}$
H 1426+428
$6.00\times 10^{-19}$ $3.09^{+0.38}_{-0.34}$ $1.58^{+0.05}_{-0.06}$ $3.16^{+1.10}_{-0.71}$
$1.20\times 10^{-18}$ $2.88^{+0.35}_{-0.37}$ $1.58^{+0.06}_{-0.07}$ $3.98^{+1.78}_{-1.03}$
$2.39\times 10^{-18}$ $2.69^{+0.40}_{-0.40}$ $1.60^{+0.06}_{-0.07}$ $5.62^{+3.92}_{-1.99}$
$4.76\times 10^{-18}$ $2.45^{+0.43}_{-0.79}$ $1.66^{+0.04}_{-0.07}$ $7.94^{+20.87}_{-3.38}$
$9.50\times 10^{-18}$ $2.24^{+0.45}_{-0.62}$ $1.73^{+0.05}_{-0.04}$ $10.23^{+21.36}_{-5.10}$
$1.89\times 10^{-17}$ $2.19^{+0.44}_{-0.49}$ $1.77^{+0.05}_{-0.03}$ $12.30^{+19.34}_{-6.27}$
$3.78\times 10^{-17}$ $2.19^{+0.38}_{-0.49}$ $1.79^{+0.05}_{-0.03}$ $12.59^{+19.03}_{-6.42}$
$7.54\times 10^{-17}$ $2.29^{+0.40}_{-0.51}$ $1.79^{+0.05}_{-0.03}$ $10.72^{+20.90}_{-5.35}$
$1.50\times 10^{-16}$ $2.40^{+0.36}_{-0.45}$ $1.79^{+0.04}_{-0.03}$ $8.71^{+13.20}_{-4.03}$
$3.00\times 10^{-16}$ $2.45^{+0.36}_{-0.41}$ $1.79^{+0.04}_{-0.03}$ $7.94^{+11.12}_{-3.48}$
1ES 1727+502
$6.00\times 10^{-19}$ $4.37^{+1.01}_{-0.82}$ $1.81^{+0.04}_{-0.04}$ $2.19^{+0.63}_{-0.45}$
$1.20\times 10^{-18}$ $4.27^{+0.98}_{-0.80}$ $1.82^{+0.04}_{-0.04}$ $2.24^{+0.71}_{-0.46}$
$2.39\times 10^{-18}$ $4.17^{+0.96}_{-0.78}$ $1.83^{+0.04}_{-0.04}$ $2.29^{+0.66}_{-0.47}$
$4.76\times 10^{-18}$ $4.27^{+0.86}_{-0.80}$ $1.83^{+0.04}_{-0.04}$ $2.29^{+0.66}_{-0.47}$
$9.50\times 10^{-18}$ $4.07^{+0.82}_{-0.76}$ $1.84^{+0.04}_{-0.04}$ $2.34^{+0.68}_{-0.48}$
$1.89\times 10^{-17}$ $4.17^{+0.84}_{-0.70}$ $1.84^{+0.04}_{-0.04}$ $2.29^{+0.66}_{-0.47}$
$3.78\times 10^{-17}$ $4.27^{+0.86}_{-0.72}$ $1.84^{+0.04}_{-0.03}$ $2.29^{+0.66}_{-0.47}$
$7.54\times 10^{-17}$ $4.37^{+0.88}_{-0.73}$ $1.84^{+0.03}_{-0.03}$ $2.24^{+0.58}_{-0.46}$
$1.50\times 10^{-16}$ $4.37^{+0.88}_{-0.73}$ $1.84^{+0.03}_{-0.03}$ $2.19^{+0.57}_{-0.45}$
$3.00\times 10^{-16}$ $4.37^{+0.88}_{-0.73}$ $1.84^{+0.03}_{-0.03}$ $2.24^{+0.58}_{-0.42}$
H 2356-309
$6.00\times 10^{-19}$ $5.50^{+0.81}_{-0.71}$ $1.41^{+0.06}_{-0.07}$ $1.02^{+0.21}_{-0.15}$
$1.20\times 10^{-18}$ $5.13^{+0.76}_{-0.66}$ $44^{+0.05}_{-0.06} 1.1$$12^{+0.23}_{-0.19}$
$2.39\times 10^{-18}$ $4.90^{+0.73}_{-0.63}$ $1.46^{+0.05}_{-0.06}$ $1.23^{+0.28}_{-0.23}$
$4.76\times 10^{-18}$ $4.37^{+0.65}_{-0.56}$ $1.50^{+0.05}_{-0.05}$ $1.45^{+0.37}_{-0.27}$
$9.50\times 10^{-18}$ $4.27^{+0.75}_{-0.55}$ $1.54^{+0.04}_{-0.05}$ $1.45^{+0.42}_{-0.30}$
$1.89\times 10^{-17}$ $3.98^{+0.59}_{-0.51}$ $1.57^{+0.04}_{-0.05}$ $1.66^{+0.58}_{-0.37}$
$3.78\times 10^{-17}$ $3.80^{+0.66}_{-0.49}$ $1.59^{+0.04}_{-0.05}$ $1.78^{+0.68}_{-0.43}$
$7.54\times 10^{-17}$ $3.63^{+0.64}_{-0.54}$ $1.61^{+0.04}_{-0.05}$ $1.86^{+0.83}_{-0.48}$
$1.50\times 10^{-16}$ $3.89^{+0.68}_{-0.58}$ $1.61^{+0.04}_{-0.05}$ $1.74^{+0.66}_{-0.42}$
$3.00\times 10^{-16}$ $3.63^{+0.64}_{-0.54}$ $1.62^{+0.04}_{-0.05}$ $1.95^{+0.87}_{-0.50}$